### v1

# This manuscript is a preprint and has not been peer reviewed | Version: July 2025

Abstract
Deep Neural Networks often suffer from a critical limitation known as Catastrophic Forgetting, where performance on past tasks degrades after learning new ones. This paper introduces a novel continual learning approach inspired by human learning strategies like Active Recall, Deliberate Practice and Spaced Repetition, named Task-Focused Consolidation with Spaced Recall (TFC-SR). TFC-SR enhances the standard experience replay with a mechanism we termed the Active Recall Probe. It is a periodic, task-aware evaluation of the model’s memory that stabilizes the representations of past knowledge. We test TFC-SR on the Split MNIST and the Split CIFAR-100 benchmarks against leading regularization-based and replay-based baselines. Our results show that TFC-SR performs significantly better than these methods. For instance, on the Split CIFAR-100, it achieves a final accuracy of 13.17% compared to standard replay’s 7.40%. We demonstrate that this advantage comes from the stabilizing effect of the probe itself, and not from the difference in replay volume. Additionally, we analyze the trade-off between memory size and performance and show that while TFC-SR performs better in memory-constrained environments, higher replay volume is still more effective when available memory is abundant. We conclude that TFC-SR is a robust and efficient approach, highlighting the importance of integrating active memory retrieval mechanisms into continual learning systems.
Keywords Continual Learning  $\cdot$ Catastrophic Forgetting  $\cdot$ Deep Neural Networks  $\cdot$ Replay-based Continual Learning
1 Introduction
Deep Neural Networks often suffer from a critical limitation known as Catastrophic Forgetting: performance on previously learnt tasks degrades as the model learns to perform newer ones [1]. This problem arises from the Stability-Plasticity dilemma [2] where a model must be flexible enough to learn newer tasks while simultaneously being stable enough to retain existing knowledge. This seriously affects the ability of a model to perform well in real life scenarios where the model must adapt to newer data and tasks incrementally over a period of time.
This paper aims to test an approach inspired by effective, synergistic learning strategies employed by humans, such as Active Recall [3], Deliberate Practice [4], and Spaced Repetition [5]. Active Recall involves effortfully retrieving information from memory. Deliberate practice involves studying a task until a certain level of proficiency is achieved. Spaced repetition refers to reviewing information at increasing intervals with the interval length determined by how well the task has been mastered. For example, If a student is learning Japanese, they may use flashcard applications like Anki to frequently test harder vocabulary (Active Recall), while words that are easier and recalled frequently are studied at progressively longer intervals (Spaced Repetition).
This paper proposes and evaluates a novel methodology to replicate these human learning methods within the continual learning framework. The primary contribution is Task-Focused Consolidation with Spaced Recall (TFC-SR), where the model learns new tasks while concurrently “practicing” previously learnt tasks using a form of experience replay.The core of our method is the Active Recall Probe, a periodic ’memory check’ where the model performs a forward pass on past experiences to evaluate its own memory state. The outcome of this memory check is then used by an adaptive spaced schedule that determines the intensity and frequency of future checks. For this initial investigation, A “naive” notion of mastery is used where mastery refers to retained performance on previously learnt tasks exceeding a predefined threshold. For the purpose of validation, TFC-SR is tested on the Split MNIST and then on the more challenging Split CIFAR-100 benchmarks and compared against standard continual learning baselines, including premier regularization-based methods like EWC [6] and SI [7] to demonstrate its effectiveness in alleviating Catastrophic Forgetting.
2 Methods
2.1 Task Protocols
For the purpose of demonstrating the effectiveness of TFC-SR, a simple CNN and a Resnet-18 [8] were tasked to work on two benchmark datasets, Split MNIST [9] and Split CIFAR-100 [10], respectively. In both cases, after the model had learnt a task, it had to classify all the classes it had seen so far. This was done to test how well the model retained its old knowledge after new information has been made available to it.
Training a simple CNN on MNIST is a relatively simple task and can often be described as being the “hello, world!” of Neural Networks. A basic implementation can be easily trained to achieve high accuracy on the dataset and thus, It was used to test the feasibility of TFC-SR in a relatively simple setting. MNIST consists of a total of 70000 grayscale, 28x28 images of handwritten digits 0-9. Those 70000 images are then split into two sets - 60000 for training and 10000 for testing. Split MNIST is a version of this dataset that is further divided into multiple classification tasks. In our case, It was divided into 5 tasks where each task involved the model learning and classifying a pair of digits. This was done to demonstrate continual learning where the model gets data sequentially instead of having all the training data available at once.
CIFAR-100 was chosen to represent a better, relatively more difficult benchmark after TFC-SR demonstrated effectiveness within a simple setting. CIFAR-100 consists of 60000 color, 32x32 images representing 100 classes. There are 600 images per class and the 100 classes are grouped into 20 different superclasses. Similar to MNIST, CIFAR-100 is also split into two sets of training and testing sets with each containing 50000 and 10000 images respectively. Since this was going to be too difficult for a simple CNN to handle, a Resnet-18 was chosen as the model of choice. For our purposes, CIFAR-100 was divided into 10 tasks, with 10 classes per task.
2.2 Model Architecture
All models were implemented using Pytorch and were trained from scratch. For the first benchmark of Split MNIST, a standard CNN was used with convolution blocks followed by a classifier head. The first convolution block consisted of a 2D convolution layer with 32 3x3 kernels followed by a ReLU activation function and a 2x2 Max Pooling layer. The second convolution block was identical to the first one with the only difference being that the 2D convolution layer had 64 3x3 kernels. The output from these blocks would then be flattened and then passed to the classifier head, which consisted of two fully connected layers. The first fully connected layer 128 units with ReLU activation while the second one was a linear layer with 10 units representing the 10 digit classes.
For Split CIFAR-100, a ResNet-18 model was utilized. To adapt the model to the CIFAR dataset’s 32x32 image size, two conventional modifications were made: The first being that the initial 7x7 convolutional layer was replaced with a 3x3 convolution layer, and the subsequent max-pooling layer was removed by replacing it with an identity mapping. The final fully connected layer was replaced with a new linear layer with 100 units to represent the 100 classes in the CIFAR-100 dataset.
2.3 Learning Algorithms and Baselines
For both Split MNIST and Split CIFAR-100, TFC-SR was compared against the following methods:
  • Sequential Fine tuning (Baseline). This represented the base, low bound benchmark without any additional method and was used to evaluate if there was any benefit to using TFC-SR at all.
  • Standard Experienced Replay [11] where the model had access to data from past tasks. Since TFC-SR is also a form of replay based method, Standard ER helped provide a better view into how the addition of Active Memory Probe helps.
  • Elastic Weight Consolidation (EWC) [6] is a premier regularization based method that adds a regularization loss in order to protect weights that are important for previous tasks based on Fisher Information Matrix.
  • Synaptic Intelligence (SI) [7] is another leading regularization based method. SI protects weights that have contributed significantly to reducing the loss function during training of past tasks. The regularization strength, $\lambda$, was found using coarse hyperparameter tuning for both EWC and SI.
2.4 TFC-SR algorithm
TFC-SR enhances the standard experience replay framework using an adaptive, scheduling mechanism based on active recall and spaced repetition. The training process for each task consists of two main components: continuous mixed batch training and an adaptive active recall schedule.
Continuous Mixed Batch training: For all tasks, the model is trained on mixed batches of data. Each batch is composed of a 50/50 split of new examples from the current task and examples randomly sampled from a replay buffer containing data from all the past tasks. To maintain a diverse and representative memory of the past experiences within a fixed memory budget, a Reservoir Replay Buffer of a fixed capacity was used.
Active Recall Probe and Adaptive Scheduling: The core novelty of TFC-SR is its adaptive scheduling which we term Active Recall Probe. This probe is a “memory check” performed periodically after a training epoch. It consists of evaluating the model’s performance on a batch of samples from the replay buffer. To get an accurate measure of retained knowledge in a class-incremental setting, this evaluation is performed in a task-aware manner, where the model’s output logits are masked to only include classes present in the replay buffer.
The outcome of the probe dictates the schedule for the next one. If the model’s performance on the replay buffer exceeds a predefined mastery_threshold, its memory is considered stable and the time until the next probe (replay_gap) is increased by a multiplicative factor. Otherwise, If the performance is below the threshold, the model’s memory is considered weak, and the next probe is scheduled for the very next epoch to encourage more intensive consolidation. The full procedure is detailed below in Algorithm 1.
2.4.1 Algorithm 1
Require: Model m, Optimizer o, Criterion c, Task Experience e
Require: Raplay Buffer b, Training DataLoader l
Require: Hyperparameters: epochs, mastery_threshold, initial_gap,
gap_multiplier
Add new data from e to b
replay_gap = initial_gap
replay_timer = replay_gap
for epoch from 1 to epochs do
m.train()
for new_data, new_targets in l do
if len(b) is sufficient then
old_data, old_targets = b.sample()
mixed_data = concat(new_data, old_data)
mixed_targets = concat(new_targets, old_targets)
o.zero_grad()
outputs = m(mixed_data)
loss = c(outputs, mixed_targets)
loss.backward()
o.step()
end if
end for
if epoch == replay_timer and task_id > 0 then
m.eval()
replay_perf = evaluate_replay_buffer(m, b)
if replay_perf >= mastery_threshold then
replay_gap = replay_gap * gap_multiplier
replay_timer = replay_timer + replay_gap
else
replay_timer = replay_timer + 1
end if
end if
end for
2.5 Code availability
The code for the experiments is available on Github: https://github.com/pritalb/tfc-sr
3 Experiments and Results
3.1 Implementation Details
All experiments were implemented using Pytorch and trained on a single NVIDIA V100 GPU. To avoid randomness and allow fair comparison between all methods, a constant seed (42) was used.
A simple CNN was used for Split MNIST while for the more challenging Split CIFAR, a modified ResNet-18 architecture was used, as detailed in section 2.2. Unless otherwise specified, all models were trained using the Adam optimizer with a learning rate of 0.001, determined via a hyperparameter tuning process on the baseline model. The batch size was 64 for both benchmarks. For Split MNIST, 10 epochs were used per task while 20 were used for Split CIFAR-100.
The key hyperparameters for each continual learning method were selected using coarse hyperparameter tuning to ensure a competitive comparison. For our proposed TFC-SR, we analyzed its sensitivity to both buffer capacity and mastery threshold (see Appendix A and B for full details). As an extension, Buffer size sensitivity was also measured for Standard ER. For the regularization based methods EWC and SI, the regularization strength, $\lambda$, was also selected via a coarse hyperparameter tuning over several orders of magnitude. The final values used for each method are listed in Table 1.
Table 1: Final Hyperparameter values for main experiments
Benchmark Method Hyperparameter Value
Split MNIST EWC Regularization $\lambda$ 10000.0
SI Regularization $\lambda$ 100.0
Replay Based (ER, TFC-SR) Buffer Capacity 200
TFC-SR Mastery Threshold 95.0
Split CIFAR-100 EWC Regularization $\lambda$ 10000.0
SI Regularization $\lambda$ 1.0
Replay Based (ER, TFC-SR) Buffer Capacity 1000
TFC-SR Mastery Threshold 10.0
All replay-based methods used a replay-to-new-data ratio of 0.5 in each batch. The TFC-SR scheduler used an initial replay gap of 1 and a gap multiplier of 1.5 for all experiments.
3.2 Evaluation on Split MNIST
To validate the feasibility of TFC-SR, experiments were conducted on the Split MNIST benchmark. Figure 1 illustrates the learning curves, plotting the average accuracy on all seen tasks after each of the five sequential tasks were completed.
The results clearly demonstrate the severity of catastrophic forgetting. After training on 5 tasks, The baseline (Sequential finetuning) model’s performance dropped to a final accuracy of 10.28%. The regularization methods provided modest improvements over the baseline with EWC scoring 20.37% accuracy while SI got a final accuracy of 32.58%.
A significant performance improvement was observed with the replay-based methods. The standard ER baseline achieved a final accuracy of 90.44%, showing that even non-adaptive replay is highly effective for this benchmark. The proposed TFC-SR method also achieved a strong and comparable final accuracy of 86.22%. Notably, the learning trajectory of TFC-SR was competitive with Standard ER throughout the experiment, briefly outperforming it until the third task before finishing at a similar level by the fourth task. The strong showing of TFC-SR motivated further analysis on a more challenging dataset.
Figure 1: Performance comparison of TFC-SR against baseline methods on the Split MNIST benchmark. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis). Higher is better.
3.3 Scalability Analysis on Split CIFAR-100
After the feasibility of the proposed TFC-SR was established, It was tested on a much harder Split CIFAR-100 benchmark in order to perform Scalability Analysis. The results are outlined in Figure 2. As expected, all methods experienced a significant drop in performance on this harder benchmark. Sequential fine tuning only managed a final performance of 7.27%. The regularization based methods, EWC and SI, failed to provide any substantial improvement and had final average accuracy of 6.16% and 8.15% respectively. Standard ER, which showed strongest results on previous MNIST benchmark, ended up with a final average accuracy of 7.4%.
In contrast, TFC-SR maintained a distinct performance advantage throughout the 10-task sequence, achieving a final performance of 13.17%. This result, which is more than 1.6 times the next best method (SI), demonstrates that TFC-SR’s active recall mechanism scales effectively to more complex and challenging continual learning scenarios.
Figure 2: Performance comparison of TFC-SR against baseline methods on the Split CIFAR-100 benchmark. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis). Higher is better.
3.4 Ablation Study: Effect of Buffer Capacity
During the initial hyperparameter tuning for TFC-SR, It was observed that the model’s performance on its memory checks was consistently high, suggesting the memory retention task was not sufficiently challenging with a large buffer. To investigate the behaviour of our method under varying degrees of memory pressure, an ablation study on the replay buffer capacity was conducted. The performance of TFC-SR against Standard ER was compared across four buffer sizes: 100, 500, 1000, and 2000. The full learning curves for each run are detailed in Appendix B, while the final accuracies are summarised in Table 2 and plotted in Figure 3. The results reveal a clear and interesting trade-off between the two methods.
In low-memory (capacity=100) and medium-memory (capacity = 500, 1000) settings, TFC-SR consistently outperforms Standard ER. The performance advantage is more at a capacity of 1000, where TFC-SR achieves a final accuracy of 13.17% against ER’s 7.40%. This suggests that when memory resources are constrained, the active recall probe and adaptive scheduling of TFC-SR provide a significant benefit over a non-adaptive replay strategy.
However, in the high-memory setting (capacity = 2000), this trend reverses. Standard ER achieves a final accuracy of 21.37%, surpassing TFC-SR’s 19.72%. This indicates that when the memory buffer is large and diverse enough, the sheer volume of brute-force replay becomes a highly effective strategy.
Table 2: Final average accuracy (%) after 10 tasks on Split CIFAR-100 for standard ER and TFC-Sr with varying replay buffer capacities.
Buffer Capacity Standard ER (%) TFC-SR(%)
100 8.51 8.55
500 9.31 9.48
1000 7.40 13.17
2000 21.37 19.72
Figure 3: Effect of buffer capacity on final model performance. The plot shows the final average accuracy on all tasks seen so far (y-axis) for Standard ER and TFC-SR at different values of buffer capacity (x-axis) on the Split CIFAR-100 benchmark. Higher is better.
3.5 Efficiency Analysis
In addition to accuracy, The computational efficiency of the replay based methods was measured using the total amount of replay batches performed during training. Additionally, For TFC-SR, the total number of Memory checks performed was also measured. The results for the primary experiments (buffer_capacity = 1000) and the “spaced replay” variant of TFC-SR are summarised in Table 3.
For the main experiments, both TFC-SR and Standard ER performed an identical number of replay batches (15800) with the only difference being the number of memory checks performed by TFC-SR (50). This indicates that the performance advantage of TFC-SR (13.17% vs 7.40%) stems from the stabilizing effect of its active recall probes, rather than the replay volume.
The trade-off between replay volume and accuracy was further explored in Appendix C, where a variant of TFC-SR with spaced replay achieved competitive performance of 10.18% - surpassing Standard ER - with over 77% fewer replay batches. This demonstrates the potential for significant computational savings when an adaptive schedule is allowed to reduce replay frequency.
Table 3: Efficiency comparison of replay based methods on Split CIFAR-100 benchmark. All methods use a buffer capacity of 1000.
Method Total Replay Batches Memory Checks Final Accuracy (%)
Standard ER 15800 0 7.40
TFC-SR 15800 50 13.17
TFC-SR (Spaced Replay) 3555 45 10.18
4 Discussion
As detailed in Section 3, TFC-SR was compared against other prominent Continual Learning methods on two benchmarks - Split MNIST and Split CIFAR-100. On both benchmarks, it demonstrated strong performance and even significantly outperformed all other methods on the difficult Split CIFAR-100 benchmark. This demonstrated that TFC-SR is a highly effective strategy for continual learning. The results from the ablation studies further revealed a key trade-off, showing that TFC-SR’s advantage is more pronounced in memory-constrained settings.
The performance gain of TFC-SR was achieved despite both methods performing an identical number of replay batches and the only difference being the presence of the active recall mechanism in TFC-SR. This suggests that the benefit stems not from the volume of replay, but from the act of memory check itself. We frame this mechanism as “Active Recall Probe”. We posit two primary reasons for the stabilizing effect. First, the repeated forward passes on past experiences force the network’s Batch Normalization layers to maintain their running statistics for older distributions, directly countering a known source of representational drift. Second, the periodic activation of neural pathways associated with past memories may act as a form of implicit regularization, strengthening those representations against interference from new learning. Additionally, This mechanism is incredibly efficient requiring minimal computational resources. As mentioned in section 3.5, It only required 50 probes for the entire Split CIFAR-100 benchmark.
TFC-SR contributes a novel mechanism to the family of replay-based continual learning methods. Unlike premier regularization methods such as EWC [6] and SI [7], which indirectly protect old knowledge by penalizing weight changes, TFC-SR directly reinforces past representations through its Active Recall Probe. Our approach also differs from other advanced replay methods. While excellent methods like iCaRL [12] focus on intelligent exemplar management for replay buffer and interesting generative methods like those in Van De Ven et al., 2020 [13] and Tadros et al., 2022 [14] focus on creating replay data without storage, TFC-SR introduces a new dimension: the process of active, periodic memory retrieval as a learning mechanism in its own right.
However, we acknowledge that this study is not without its own limitations. First, our experiments were conducted exclusively on image classification benchmarks; the efficacy of TFC-SR on other data modalities such as Natural Language Processing or Reinforcement Learning tasks remain to be explored. Second, our replay buffer used a standard reservoir sampling strategy. Incorporating more advanced buffer management techniques, such as iCaRL’s herding algorithm, Iterative Projection and Matching (IPM) [15] or gradient-based sample selection [16] could provide further improvements. Lastly, We used a naive notion of “mastery” for the adaptive scheduler. It was based on a simple performance threshold. A more nuanced metric could unlock more complex and efficient scheduling dynamics.
The findings from this paper open several exciting avenues for future research. Our preliminary results with a ’Spaced Replay’ variant (Appendix C) show a promising trade-off between accuracy and computational cost, warranting a deeper investigation into sparse, episodic consolidation. Furthermore, the viability of an alternative ’Mastery-Gated Progression’ paradigm (Appendix D) suggests a rich area for future research into self-directed and adaptive learning schedules. Ultimately, this work strongly supports the hypothesis that building computational analogues of human cognitive processes, like active recall, is a fruitful direction for developing more robust, efficient, and capable artificial intelligence systems.
Acknowledgments
I would like to offer my genuine gratitude to all the people who have provided their direct or indirect support throughout the completion of this work. I also acknowledge the significant contribution of AI assistance in improving readability and grammatical structure of this paper as well as in saving significant setup time during experimentation. I also would like to extend my thanks to my family and friends for their unwavering support and encouragement.
Appendix A TFC-SR’s Sensitivity to Mastery Threshold
To find the optimal mastery threshold for TFC-SR, we tested its performance across a range of values on the CIFAR-100 benchmark: (10.0, 20.0, 30.0, 50.0, 70.0, 90.0, 99.0). We found that performance was robust across all values. The results are outlined in Table A.1 and illustrated in Figure A.1.
For thresholds between 10% and 90%, the final accuracy was nearly identical (13.17%), indicating that the model’s performance on the replay buffer consistently exceeded the targets. This caused the adaptive scheduler to adopt its most efficient mode, performing the minimum number of memory checks (50-63).
The 99% threshold served as a crucial stress test, creating a condition where the memory check would only succeed if the model’s retention on past knowledge was nearly perfect. The scheduler performed 103 memory checks in this configuration, significantly more than in the lower-threshold runs. This confirms that the check was frequently failing, forcing more intensive probing as designed.
However, the fact that only 103 checks were performed, which is close to only half of the maximum possible, proves that the model did succeed in meeting the 99% threshold on multiple occasions. This indicates moments of powerful, near-perfect memory consolidation even while learning new tasks.
Most notably, the final accuracy of this 99% threshold run was 11.58%. This result significantly outperforms the Standard ER baseline (7.40%), demonstrating the powerful benefit of Active Recall Probe mechanism itself. Additionally, this performance is only slightly lower than the performance of the best-performing threshold values (13.17%), showing that the algorithm is robust and does not catastrophically fail even when subjected to a highly demanding and volatile replay schedule.
Table A.1: TFC-SR’s performance across multiple mastery thresholds on CIFAR-100 benchmark
Mastery Threshold Memory Checks Final Accuracy (%)
10.0 50 13.17
20.0 50 13.17
30.0 50 13.17
50.0 50 13.17
70.0 51 12.45
90.0 63 13.17
99.0 103 11.58
Figure A.1: Learning curves of TFC-SR with varying mastery thresholds on Split CIFAR-100. For comparison, the performance of Standard ER baseline (buffer capacity = 1000) is included. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis)
Appendix B Full Learning Curves for Buffer Capacity Analysis
This section provides the detailed learning curves that support the ablation study on replay buffer capacity presented in Section 3.4.
We evaluated both TFC-SR and Standard ER across four different buffer capacities: 100, 500, 1000, and 2000. The full, task-by-task learning trajectory for each TFC-SR run is shown in Figure B.1. The corresponding trajectories for Standard ER baseline are shown in Figure B.2. The final accuracy values from these curves were used to generate the trend plot in Figure 3 and the summary in Table 2 of the main paper.
Figure B.1: Learning curves of TFC-SR with varying buffer capacities on Split CIFAR-100. Each line represents a different buffer capacity (100, 500, 1000, 2000). The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis)
Figure B.2: Learning curves of Standard ER with varying buffer capacities on Split CIFAR-100. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis)
Appendix C Analysis of a Spaced Replay Variant (TFC-SR²)
To further investigate the trade-off between replay volume and final performance, we implemented and tested a “Spaced Replay” variant of TFC-SR, which we informally named TFC-SR2. In this setup, the continuous mixed batch replay is disabled; instead the replay is performed only during the specific epochs that are triggered by the adaptive scheduling mechanism. On all other epochs, the model trains exclusively on data from the new task. This experiment allows us to isolate the impact of the scheduled replay events themselves, under a drastically reduced computational budget. The results of this experiment are compared against our primary TFC-SR model and the Standard ER baseline in Table 3 of the main paper while the full learning curves are illustrated in Figure C.1.
The TFC-SR (Spaced Replay) variant achieved a final accuracy of 10.18%. This result is highly significant for two reasons. First, despite using over 77% fewer replay batches (3555 vs 15800), it substantially outperforms the brute-force Standard ER baseline (7.40%), demonstrating that even a small amount of intelligently scheduled replay is more effective than a high volume non-adaptive replay. Second, while its accuracy is lower than our main continuous replay variant (13.17%), this result confirms that a high volume of replay is still beneficial for achieving maximum performance on a challenging benchmark. This finding highlights a clear and practical trade-off between computational efficiency and performance.
Figure C.1: Learning curves comparing TFC-SR (Spaced Replay) against our primary TFC-SR model and the Standard ER baseline on Split CIFAR-100. All methods used a buffer size of 1000. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis).
Appendix D Mastery-Gated Progression
In developing our approach, we explored a second, more literal interpretation of the “deliberate practice” analogy, which we named Mastery-Gated Progression (MGP). Unlike TFC-SR, which interleaves learning and replay, MGP separates them into two distinct phases. In this paradigm, the model trains on a new task until it achieves a performance greater than a predefined mastery threshold. Only after this criterion is met does the model “progress” to the next task. This creates an adaptive training schedule where the time spent on each task is an emergent property of the task’s difficulty.
The mastery criterion is two-fold: the model’s accuracy on the current task’s test set must exceed a new_task_mastery_thresh, and its performance on past tasks (measured via evaluate_replay_buffer) must remain above a retention_thresh. A max_epochs_per_task hyperparameter is used as a safeguard to prevent infinite training loops on exceptionally difficult tasks. The full logic is detailed in Algorithm D.1.
We tested MGP on the Split CIFAR-100 benchmark. Its performance is compared against the primary baselines in Figure D.1. The MGP model achieved a final accuracy of 11.63%, a result that is highly competitive with the Standard ER baseline (7.40%) and demonstrates the viability of this alternative learning paradigm.
The most insightful finding from this experiment was the variable number of epochs required to achieve mastery for each task: (47, 9,, 22, 50, 50, 9, 15, 18, 7, 19). This result makes the inherent difficulty of each incremental task explicit. For example, the model required 47 epochs to learn the first task from a random initialization, but only 9 for the second, suggesting a degree of positive knowledge transfer. This dynamic allocation of computational effort stands in contrast to fixed-epoch schedules and presents a rich area for future research into self-directed learning [17].
Algorithm D.1 Mastery Gated Progression (MGP) Training for a single task
Require: Model m, Optimizer o, Criterion c, Task Experience e
Require: Replay Buffer b, task_id, Dataloader l
Require: Hyperparameters: max_epochs, mastery_threshold, retention_threshold
epoch = 0
mastery_achieved = false
while not mastery_achieved and epoch < max_epochs do
epoch = epoch + 1
m.train()
for new_data, new_targets in l do
if task_id > 0 then
// perform mixed batch training
old_data, old_targets = b.sample()
mixed_data = concat(new_data, old_data)
TrainStep(m, o, c, mixed_data, mixed_targets)
else
// perform new-task-only training
TrainStep(m, o, c, new_data, new_targets)
end if
end for
// perform mastery check
m.eval()
new_task_perf = evaluate(m. e.test_dataset)
retention_perf = evaluate_replay_buffer(m,, b)
if new_task_perf >= mastery_thresh and retention_perf >= retention_thresh then
mastery_achieved = True
end if
end while
return epoch
Figure D.1: Learning curves comparing MGP against TFC-SR and the Standard ER on Split CIFAR-100. All methods used a buffer size of 1000. The plot shows the average accuracy on all tasks seen so far (y-axis) after each new task is learnt (x-axis).